# Understanding LLM Agent Behaviours via Game Theory: Strategy Recognition, Biases and Multi-Agent Dynamics

### v1

@@ Abstract
As Large Language Models (LLMs) increasingly operate as autonomous decision-makers in interactive and multi-agent systems and human societies, understanding their strategic behaviour has profound implications for safety, coordination, and the design of AI-driven social and economic infrastructures. Assessing such behaviour requires methods that capture not only what LLMs output, but the underlying intentions that guide their decisions. In this work, we extend the FAIRGAME framework to systematically evaluate LLM behaviour in repeated social dilemmas through two complementary advances: a payoff-scaled Prisoner’s Dilemma isolating sensitivity to incentive magnitude, and an integrated multi-agent Public Goods Game with dynamic payoffs and multi-agent histories. These environments reveal consistent behavioural signatures across models and languages, including incentive-sensitive cooperation, cross-linguistic divergence and end-game alignment toward defection. To interpret these patterns, we train traditional supervised classification models on canonical repeated-game strategies and apply them to FAIRGAME trajectories, showing that LLMs exhibit systematic, model- and language-dependent behavioural intentions, with linguistic framing at times exerting effects as strong as architectural differences. Together, these findings provide a unified methodological foundation for auditing LLMs as strategic agents and reveal systematic cooperation biases with direct implications for AI governance, collective decision-making, and the design of safe multi-agent systems.
@@ 1 Introduction
Large language models (LLMs) are no longer restricted to passively generating text, but they now support a wide range of applications. They are increasingly deployed as agents that interact both with human users and with other agents in recommendation systems, negotiation tools, and multi-agent assistants [tessler2024ai, patel2020leveraging, hammond2025multi]. In these settings, LLMs repeatedly face cooperation dilemmas: they must decide whether to contribute to a shared goal, free-ride on others’ efforts, or enforce social norms [lu2024llms, tessler2024ai]. Evaluating such systems therefore requires more than checking whether individual answers are factually correct or polite; we also need to understand the behavioural intentions or strategies that LLM agents implement over time, and how these depend on incentives, language, and role [buscemi2025llms, akata2025, jia2025large, sun2025game, mao2025alympics].
We adopt the notion of “behavioural intention” from the literature on repeated games, where an intention is operationalised as a decision rule that maps past interaction histories to current actions. Classical game theory has long studied how such strategies emerge, stabilise, and can be inferred in repeated social dilemmas. Recent machine learning work builds on this by showing that canonical strategies—such as Always Cooperate (ALLC), Always Defect (ALLD), Tit-for-Tat (TFT), and Win-Stay-Lose-Shift (WSLS)—can be recognised in the Iterated Prisoner’s Dilemma (IPD) by training classifiers on noisy trajectories of play between artificial agents [diStefanoIntention2023]. This line of research suggests that behavioural intention can be treated as a learnable object: given enough action trajectories, we may infer which strategy best explains an agent’s behaviour, even in the presence of execution noise.
@@
In this work, we bring these strands together and extend FAIRGAME along two complementary directions. From a game-theoretic perspective, we design two families of repeated-game experiments for LLM agents. First, we introduce a payoff-scaling module for the Prisoner’s Dilemma that multiplies all entries of a fixed payoff matrix by a scalar factor, thereby manipulating the stakes of cooperation while preserving the underlying strategic structure. This allows us to ask whether LLM agents are sensitive to the absolute magnitude of payoffs, and whether such sensitivity depends on the model family or the language of interaction. Second, we develop a genuinely multi-agent extension based on a three-player Public Goods Game (PGG) [sigmund:2010bo] with configurable multiplication factors, languages, and injected personalities. This multi-agent module enables us to study free-riding, coordination, and coalition-like behaviour in collective social dilemmas where payoffs depend on group-level contributions rather than pairwise interactions.
From a data-driven perspective, we then treat the trajectories generated by these experiments as input to a machine learning pipeline for behavioural intention recognition. Following the protocol of [diStefanoIntention2023], we generate synthetic IPD trajectories for the four canonical strategies (ALLC, ALLD, TFT, WSLS) under execution noise and train a suite of classifiers-including Logistic Regression [logit], Random Forests, and Long Short-Term Memory (LSTM) networks [lstm]-to identify which architectures are most robust to noisy play. The best-performing model is then used as an intent recogniser on FAIRGAME logs: we encode LLM game histories into state-action sequences and infer which canonical strategies-and which mixtures thereof-best explain the observed behaviour. This allows us to move beyond raw cooperation rates and examine how LLM agents’ latent intentions vary across models, languages, personalities, and roles (e.g., first-mover vs. second-mover).
@@
  • RQ3. To what extent can we predict and classify the behavioural intentions of LLM agents in repeated cooperation dilemmas using machine learning, and what systematic biases-across models, languages, personalities, and positional roles-emerge when we interpret their gameplay through canonical strategy classes?
@@
2 Methodology
2.1 FAIRGAME: Extension
2.1.1 FAIRGAME overview
FAIRGAME (Framework for AI Agents Bias Recognition using Game Theory) [buscemi2025fairgame] provides the foundational computational infrastructure upon which our work is built. The framework was originally introduced to support systematic, reproducible evaluations of Large Language Models (LLMs) through controlled multi-agent game-theoretic experiments. It offers a unified pipeline for defining games, orchestrating interactions among LLM agents, and analysing emergent behavioural patterns across languages, personalities, and strategic configurations.
@@ 2.1.1 FAIRGAME overview
In its original formulation, FAIRGAME has been primarily applied to symmetric two-player matrix games, such as variants of the Prisoner’s Dilemma and coordination games, to compare the behavior of different LLMs and languages under fixed payoff structures. In this work, we extend FAIRGAME along two complementary directions. First, we introduce a payoff-scaling module for the Prisoner’s Dilemma to investigate how sensitive LLM agents are to changes in the magnitude of incentives. Second, we develop a multi-agent extension that instantiates a multi-player Public Goods Game, generating richer trajectories of group interaction that we subsequently analyze using machine learning methods for strategy and intent recognition.
2.1.2 Game stakes: Prisoner’s Dilemma payoff-scaling
We first examine the sensitivity of LLM agents to the absolute magnitude of incentives in a dyadic setting. To this end, we use a repeated Prisoner’s Dilemma in which only the numerical values of the payoffs are scaled, while the underlying strategic structure of the game is kept fixed. In this way, the “stakes” of the interaction are varied without changing best responses or the ranking of outcomes.
@@ 2.1.2 Game stakes: Prisoner’s Dilemma payoff-scaling
Two-player games between LLM agents are run using FAIRGAME as the simulation engine. Each game is played for a fixed, finite horizon of $T=10$ rounds, and in every round both agents observe the full public history of past actions and payoffs before choosing their next move. Agents do not communicate outside of their action choices. For each parameter configuration, we simulate multiple independent runs to account for the stochasticity of LLM outputs. We evaluate three LLM backends: GPT-4o [gpt4o], Claude 3.5 Haiku [claude3], and Mistral Large [mistral]. To probe potential cross-lingual effects, the same game is instantiated in English and Vietnamese. In all conditions, neutral framing is employed: “Option A” corresponds to defection and “Option B” to cooperation, and the prompt does not contain any explicit moral or normative language. We tested and evaluated the results of 40 games with 400 decisions per setting.
2.1.3 Multi-agent extension: Public Goods Game
@@
which induces the standard public-goods social dilemma: collective welfare is maximised when all agents contribute, but unilateral defection strictly increases individual payoff and enables free-riding.
To support this group interaction within FAIRGAME, the original two-player matrix game implementation is extended along three axes. First, the static $2\times 2$ payoff matrix is replaced by a dynamic public-goods payoff module, which computes the vector of payoffs $\boldsymbol{\pi}_{t}$ from the joint action profile $\mathbf{s}_{t}$ according to Eq. (1). Second, the game history is generalised from bilateral outcomes to vector-valued records: for each round, FAIRGAME now stores the full tuple of actions and payoffs for all $N$ agents, enabling strategy analysis at the group level. Third, the prompt-generation mechanism is adapted so that LLM agents reason about multi-agent histories and group-level incentives rather than dyadic exchanges. Public-goods-specific templates present the rules, worked payoff examples derived from $(c,r,N)$, the current round index, and the full multi-agent history, and are instantiated separately for each agent.
The resulting execution loop preserves FAIRGAME’s overall control flow but operates over joint strategy profiles and dynamic payoffs. Algorithm 1 summarises this extension of FAIRGAME Algorithms 2-3.
@@
    // dynamic public-goods payoff
@@
Here QueryAllAgents generalises FAIRGAME’s per-round interaction to the multi-agent PGG setting: for each agent in $G$, a public-goods-specific prompt is constructed using the current round index, game parameters $(c,r,N)$, and the full multi-agent history; the underlying LLM is queried; and the textual response is parsed into a discrete action in $\{Contribute,Keep\}$. The procedure PGGPayoff applies Eq. (1) to the resulting joint strategy profile $\mathbf{s}_{t}$ to produce the payoff vector $\boldsymbol{\pi}_{t}$. Full pseudocode for game instantiation, the execution loop, and the per-round decision routine is provided in Appendix A.2.
The experimental configuration systematically varies the incentive structure, language, and model family. Games are played for a fixed horizon of $T=10$ rounds, and agents are informed of this horizon in advance. For each condition, three LLM-based agents (one per player) are instantiated, the public-goods payoff module defined by Eq. (1) is attached, and prompts are generated via the PGG template described above. In every round, all agents are queried with their respective prompts, their textual responses are parsed into actions in $\{\text{Contribute},\text{Keep}\}$, and the resulting joint action profile and payoff vector are appended to the game history.
@@
2.2 Machine Learning Approaches for Understanding Behavior LLMs
@@
Figure 1: Machine Learning Pipeline for Understanding Behaviour LLMs. Starting from action sequences associated with canonical strategies (ALLC, ALLD, TFT, GTFT, WSLS) under varying noise conditions, we train machine learning models to infer and classify underlying behavioural intentions. We then apply the best-performing model to the gameplay data generated by FAIRGAME. High-confidence predictions are used to identify which strategies the LLM adopts, whereas low-confidence cases are reserved for subsequent analysis to investigate the possibility of emerging behaviours in the LLM.
@@ 2.2.1 Training Phase
Following the protocol and the synthetic interaction trajectories based on four canonical strategies: Tit-for-Tat (TFT), Always Cooperate (ALLC), Always Defect (ALLD), and Win-Stay-Lose-Shift (WSLS) defined in [diStefanoIntention2023]. Crucially, to accommodate the stochastic imperfections inherent in generative AI, a noise parameter ($\epsilon\in\{0,0.05\}$) is injected into these sequences. This robust synthetic dataset is then used to train and benchmark a suite of classifier architectures, including Logistic Regression [cox1958regression], Random Forests [rf], Neural Networks [nn], and Long Short-Term Memory (LSTM) [lstm] networks-allowing us to identify the optimal model for capturing temporal strategic dependencies. The trained model is subsequently used to predict the strategies underlying the gameplay turns of LLM agents.
@@ 2.2.2 Inference Phase
The FAIRGAME simulation logs record agent interactions using neutral labels: OptionA (representing Defection) and OptionB (representing Cooperation). Our encoding scheme transforms these raw sequences into a semantic "state-action" format that explicitly captures the conditional nature of game-theoretic strategies. We then determine the interaction outcome for each round t1t-1 based on the joint actions: Reward ($R$, mutual cooperation), Punishment ($P$, mutual defection), Temptation ($T$, agent defects while opponent cooperates), and Sucker ($S$, agent cooperates while opponent defects).
The chosen model is applied to the processed FAIRGAME trajectories. For each agent $i$ in a game session, the model outputs a probability distribution over the four strategy classes. To rigorously analyze the results and avoid over-interpreting ambiguous behaviors, we focus our downstream analysis on high-confidence predictions, defined as classifications where the model assigns a probability greater than 0.9 to a single strategy. Predictions with confidence below 0.9 require separate analysis to determine whether they reflect emerging LLM behaviors that differ from those of humans.
@@ 3.1 Payoff Magnitude Sensitivity in Prisoner’s Dilemma
Because the payoff matrix is scaled by $\lambda$, the range of total penalties scales accordingly. Figure 2 shows that the very low–magnitude payoff setting (i.e., $\lambda=0.1$) consistently produces higher overall penalties across models, languages, and personality pairings. This indicates that reducing the payoff scale leads to more frequent defections. Within each language, the ordinary and high payoff settings yield broadly similar patterns, with only minor variations at specific points.
When comparing languages, the results indicate that several LLMs are sensitive to linguistic context. Cooperative pairings in Vietnamese context strongly favour defection, considering the two agents GPT-4o and Mistral. Notably, when shifting from English to Vietnamese, LLM agents often reverse their behaviour: models that yield lower total penalties in English tend to produce higher penalties in Vietnamese, and vice versa.
@@
Figure 3 presents the strategy trajectories across rounds for each model under varying values of the scaling parameter $\lambda$. For Claude 3.5 Haiku, the relationship between payoff magnitude and strategic behaviour appears relatively weak and inconsistent. GPT-4o, however, shows a clearer pattern, exhibiting increasingly selfish behaviour as the payoff matrix is scaled down. In contrast, Mistral Large demonstrates the opposite tendency, with its behaviour shifting counter to the trend observed in GPT-4o.
The performance trajectories of the Claude 3.5 Haiku model in Figure 3 exhibit a general downward trend as the number of iterations increases. The very low trajectory, predominantly situated in the upper quadrant (favoring Option A), displays a gradual shift towards the lower quadrant (Option B), interrupted by a minor peak at Round 5 that suggests a notable degree of strategic volatility. Conversely, the high and ordinary trajectories are primarily concentrated in the lower domain, indicating a cooperative tendency designed to yield a higher average expected utility. The very low line demonstrates a significantly greater inclination towards the upper spectrum compared to the high line, a divergence that may be partly attributed to linguistic bias. Specifically, within the context of the Vietnamese language combined with a payoff scalar of $\lambda=0.1$, Claude 3.5 Haiku appears to misinterpret the unit magnitude of the payoff matrix, resulting in the erroneous maximization of penalties rather than their intended minimization.
@@
Regarding the Mistral Large model, empirical observations indicate that higher payoff values correlate with an increased likelihood of defection, a behaviour consistent with the dominant strategy concept in game theory. Furthermore, a distinct visual correlation is evident, characterized by a substantial spike at Round 2. This anomaly suggests a retaliatory mechanism triggered by a defection in the initial round, or alternatively, represents a probing tactic aimed at minimizing the maximum potential penalty.
3.2 Public Goods Game Behavior
We now turn to the multi-agent Public Goods Game to address RQ2: how do LLM agents behave in a genuinely collective social dilemma, and how does this behavior depend on incentives, time, and framing (model, personality, language).

### v2

@@ Abstract
As Large Language Models (LLMs) increasingly operate as autonomous decision-makers in interactive and multi-agent systems and human societies, understanding their strategic behaviour has profound implications for safety, coordination, and the design of AI-driven social and economic infrastructures. Assessing such behaviour requires methods that capture not only what LLMs output, but the underlying intentions that guide their decisions. In this work, we extend the FAIRGAME framework to systematically evaluate LLM behaviour in repeated social dilemmas through two complementary advances: a payoff-scaled Prisoner’s Dilemma isolating sensitivity to incentive magnitude, and an integrated multi-agent Public Goods Game with dynamic payoffs and multi-agent histories. These environments reveal consistent behavioural signatures across models and languages, including incentive-sensitive cooperation, cross-linguistic divergence and end-game alignment toward defection. To interpret these patterns, we train traditional supervised classification models on canonical repeated-game strategies and apply them to LLM decisions in FAIRGAME, showing that LLMs exhibit systematic, model- and language-dependent behavioural intentions, with linguistic framing at times exerting effects as strong as architectural differences. Together, these findings provide a unified methodological foundation for auditing LLMs as strategic agents and reveal systematic cooperation biases with direct implications for AI governance, collective decision-making, and the design of safe multi-agent systems.
@@ 1 Introduction
Large language models (LLMs) are increasingly deployed as agents that interact with human users and with one another in recommendation systems, negotiation tools, and multi-agent assistants [tessler2024ai, patel2020leveraging, hammond2025multi]. In these settings, LLMs are repeatedly exposed to cooperation dilemmas, where they may produce behaviours that resemble contributing to a shared goal, free-riding on others, or enforcing social norms [lu2024llms, tessler2024ai, balabanova2025media, commitment]. Such behaviour does not reflect genuine intentions or internal goals. It is the result of statistical patterns learned during training, combined with the incentives and context provided during interaction. Evaluating these systems therefore requires more than verifying factual correctness or conversational quality. It requires analysing the emergent strategies that LLM-based agents tend to exhibit over time, how these strategies are shaped by prompting, reward structures, and role assignment, and how they differ across languages and tasks [buscemi2025llms, correia2025alife, akata2025, jia2025large, sun2025game, mao2025alympics, proverbio2025can, buscemi2025strategic].
We adopt the notion of “behavioural intention” from the literature on repeated games, where an intention is operationalised as a decision rule that maps past interaction histories to current actions. Classical and evolutionary game theory has long studied how such strategies emerge, stabilise, and can be inferred in repeated social dilemmas [diStefanoIntention2023, Han2012ALIFEjournal]. Recent work builds on this by showing that canonical strategies—such as Always Cooperate (ALLC), Always Defect (ALLD), Tit-for-Tat (TFT), and Win-Stay-Lose-Shift (WSLS) [Axelrod1980Effective, sigmund:2010bo]—can be recognised in the Iterated Prisoner’s Dilemma (IPD) by training classifiers on noisy trajectories of play between artificial agents [diStefanoIntention2023]. This line of research suggests that behavioural intention can be treated as a learnable object: given enough action trajectories, we may infer which strategy best explains an agent’s behaviour, even in the presence of execution noise.
@@
In this work, we bring these strands together and extend FAIRGAME along two complementary directions. From a game-theoretic perspective, we design two families of repeated-game experiments for LLM agents. First, we introduce a payoff-scaling module for the Prisoner’s Dilemma that multiplies all entries of a fixed payoff matrix by a scalar factor, thereby manipulating the stakes of cooperation while preserving the underlying strategic structure [han_when_2021]. This allows us to ask whether LLM agents are sensitive to the absolute magnitude of payoffs, and whether such sensitivity depends on the model family or the language of interaction. Second, we develop a genuinely multi-agent extension based on a three-player Public Goods Game (PGG) [sigmund:2010bo] with configurable multiplication factors, languages, and injected personalities. This multi-agent module enables us to study free-riding, coordination, and coalition-like behaviour in collective social dilemmas where payoffs depend on group-level contributions rather than pairwise interactions.
From a data-driven perspective, we then treat the trajectories (i.e. sequences of actions by LLMs) generated by these experiments as input to a machine learning pipeline for behavioural intention recognition. Following the protocol of Di Stefano et al. [diStefanoIntention2023], we generate synthetic IPD trajectories for the four canonical strategies (ALLC, ALLD, TFT, WSLS) under execution noise and train a suite of classifiers, including Logistic Regression [logit], Random Forests, and Long Short-Term Memory (LSTM) networks [lstm], to identify which architectures are most robust to noisy play. The best-performing model is then used as an intent recogniser on FAIRGAME logs: we encode LLM game histories into state-action sequences and infer which canonical strategies-and which mixtures thereof-best explain the observed behaviour. This allows us to move beyond raw cooperation rates and examine how LLM agents’ latent intentions vary across models, languages, personalities, and roles (e.g., first-mover vs. second-mover).
@@
  • RQ3. To what extent can we predict and classify the behavioural intentions of LLM agents in repeated cooperation dilemmas using machine learning, and what systematic biases (across models, languages, personalities, and positional roles) emerge when we interpret their gameplay through canonical strategy classes?
@@
2 Background
2.1 LLMs and Game Theory
2.2 FAIRGAME
FAIRGAME (Framework for AI Agents Bias Recognition using Game Theory) [buscemi2025fairgame] provides the foundational computational infrastructure upon which our work is built. The framework was originally introduced to support systematic, reproducible evaluations of LLMs through controlled multi-agent game-theoretic experiments. It offers a unified pipeline for defining games, orchestrating interactions among LLM agents, and analysing emergent behavioural patterns across languages, personalities, and strategic configurations.
@@ 2.2 FAIRGAME
In its original formulation, FAIRGAME has been primarily applied to symmetric two-player matrix games, such as variants of the Prisoner’s Dilemma and coordination games, to compare the behaviour of different LLMs and languages under fixed payoff structures. In this work, we extend FAIRGAME along two complementary directions. First, we introduce a payoff-scaling module for the Prisoner’s Dilemma to investigate how sensitive LLM agents are to changes in the magnitude of incentives. Second, we develop a multi-agent extension that instantiates a multi-player Public Goods Game, generating richer trajectories of group interaction that we subsequently analyse using machine learning methods for strategy and intent recognition.
3 Methodology
In this section we describe the methodology adopted to address the three research questions introduced in 1.
3.1 Game stakes: Prisoner’s Dilemma payoff-scaling
We first examine the sensitivity of LLM agents to the absolute magnitude of incentives in a dyadic setting. To this end, we use a repeated Prisoner’s Dilemma in which only the numerical values of the payoffs are scaled, while the underlying strategic structure of the game is kept fixed. In this way, the “stakes” of the interaction are varied without changing best responses or the ranking of outcomes. It has been shown that, more cooperative behaviours are strongly influenced by this factor in the context of repeated games [han_when_2021].
@@ 3.1 Game stakes: Prisoner’s Dilemma payoff-scaling
Two-player games between LLM agents are run using FAIRGAME as the simulation engine. Each game is played for a fixed, finite horizon of $T=10$ rounds, and in every round both agents observe the full public history of past actions and payoffs before choosing their next move. Agents do not communicate outside of their action choices. For each parameter configuration, we simulate multiple independent runs to account for the stochasticity of LLM outputs. We evaluate three LLM backends: GPT-4o [gpt4o], Claude 3.5 Haiku [claude3], and Mistral Large [mistral]. To probe potential cross-lingual effects, the same game is instantiated in English and Vietnamese. In all conditions, neutral framing is employed: “Option A” corresponds to defection and “Option B” to cooperation, and the prompt does not contain any explicit moral or normative language. We test and evaluate the results of 40 games with 400 decisions per setting.
3.1.1 Public Goods Game
@@
which induces the standard public goods social dilemma: collective welfare is maximised when all agents contribute, but unilateral defection strictly increases individual payoff and enables free-riding.
To support this group interaction within FAIRGAME, the original two-player matrix game implementation is extended along three axes. First, the static $2\times 2$ payoff matrix is replaced by a dynamic public goods payoff module, which computes the vector of payoffs $\boldsymbol{\pi}_{t}$ from the joint action profile $\mathbf{s}_{t}$ according to Eq. (1). Second, the game history is generalised from bilateral outcomes to vector-valued records: for each round, FAIRGAME now stores the full tuple of actions and payoffs for all $N$ agents, enabling strategy analysis at the group level. Third, the prompt-generation mechanism is adapted so that LLM agents reason about multi-agent histories and group-level incentives rather than dyadic exchanges. public goods-specific templates present the rules, worked payoff examples derived from $(c,r,N)$, the current round index, and the full multi-agent history, and are instantiated separately for each agent.
The resulting execution loop preserves FAIRGAME’s overall control flow but operates over joint strategy profiles and dynamic payoffs. Algorithm 1 summarises this extension of FAIRGAME.
@@
    // dynamic public goods payoff
@@
Here QueryAllAgents generalises FAIRGAME’s per-round interaction to the multi-agent PGG setting: for each agent in $G$, a PGG-specific prompt is constructed using the current round index, game parameters $(c,r,N)$, and the full multi-agent history; the underlying LLM is queried; and the textual response is parsed into a discrete action in $\{Contribute,Keep\}$. The procedure PGGPayoff applies Eq. (1) to the resulting joint strategy profile $\mathbf{s}_{t}$ to produce the payoff vector $\boldsymbol{\pi}_{t}$. Full pseudocode for game instantiation, the execution loop, and the per-round decision routine is provided in Appendix A.2.
The experimental configuration systematically varies the incentive structure, language, and model family. Games are played for a fixed horizon of $T=10$ rounds, and agents are informed of this horizon in advance. For each condition, three LLM-based agents (one per player) are instantiated, the public goods payoff module defined by Eq. (1) is attached, and prompts are generated via the PGG template described above. In every round, all agents are queried with their respective prompts, their textual responses are parsed into actions in $\{\text{Contribute},\text{Keep}\}$, and the resulting joint action profile and payoff vector are appended to the game history.
@@
3.2 Machine Learning approaches for understanding LLM behaviour
@@
Figure 1: Machine Learning Pipeline for Understanding LLM Behaviour. Starting from action sequences associated with canonical strategies (ALLC, ALLD, TFT, GTFT, WSLS) under varying noise conditions, we train machine learning models to infer and classify underlying behavioural intentions. We then apply the best-performing model to the gameplay data generated by FAIRGAME. High-confidence predictions are used to identify which strategies the LLM adopts, whereas low-confidence cases are reserved for subsequent analysis to investigate the possibility of emerging behaviours by the LLM.
@@ 3.2.1 Training Phase
Following the protocol and the synthetic interaction trajectories based on four canonical strategies: Tit-for-Tat (TFT), Always Cooperate (ALLC), Always Defect (ALLD), and Win-Stay-Lose-Shift (WSLS) defined in [diStefanoIntention2023, sigmund:2010bo]. Crucially, to accommodate the stochastic imperfections inherent in generative AI, a noise parameter ($\epsilon\in\{0,0.05\}$) is injected into these sequences. This robust synthetic dataset is then used to train and benchmark a suite of classifier architectures, including Logistic Regression [cox1958regression], Random Forests [rf], Neural Networks [nn], and Long Short-Term Memory (LSTM) [lstm] networks. That allows us to identify the optimal model for capturing temporal strategic dependencies. The trained model is subsequently used to predict the strategies underlying the gameplay turns of LLM agents.
@@ 3.2.2 Inference Phase
The FAIRGAME simulation logs record agent interactions using neutral labels: OptionA (representing Defection) and OptionB (representing Cooperation). Our encoding scheme transforms these raw sequences into a semantic "state-action" format that explicitly captures the conditional nature of game-theoretic strategies. We then determine the interaction outcome for each round (t1)(t-1) based on the joint actions: Reward ($R$, mutual cooperation), Punishment ($P$, mutual defection), Temptation ($T$, agent defects while opponent cooperates), and Sucker ($S$, agent cooperates while opponent defects).
The chosen model is applied to the processed FAIRGAME trajectories. For each agent $i$ in a game session, the model outputs a probability distribution over the four strategy classes. To rigorously analyse the results and avoid over-interpreting ambiguous behaviours, we focus our downstream analysis on high-confidence predictions, defined as classifications where the model assigns a probability greater than 0.9 to a single strategy. Predictions with confidence below 0.9 require separate analysis to determine whether they reflect emerging LLM behaviours that differ from those of humans.
@@ 4 Results
In this section, we present the results obtained using the methodology described in 3 to address the research questions targeted in this work.
@@ 4.1 Payoff Magnitude Sensitivity in Prisoner’s Dilemma
Because the payoff matrix is scaled by $\lambda$, the range of total penalties scales accordingly. Figure 2 shows that the very low–magnitude payoff setting (i.e., $\lambda=0.1$) consistently produces higher overall penalties across models, languages, and personality pairings. This indicates that when the stake of the game is very low, defection is highly frequent. This is in line with the game theory analysis in [han_when_2021] (see Figure 4 therein). Within each language, the ordinary and high payoff settings yield broadly similar patterns, with only minor variations at specific points.
When comparing languages, the results indicate that several LLMs are sensitive to the linguistic context. Cooperative pairings in the Vietnamese context strongly favour defection, considering the two agents GPT-4o and Mistral. Notably, when shifting from English to Vietnamese, LLM agents often reverse their behaviour: models that yield lower total penalties in English tend to produce higher penalties in Vietnamese, and vice versa.
@@
Figure 3 presents the sequences of choices across rounds for each model under varying values of the scaling parameter $\lambda$. For Claude 3.5 Haiku, the relationship between payoff magnitude and strategic behaviour appears relatively weak and inconsistent [razeghi2022impact]. GPT-4o, however, shows a clearer pattern, exhibiting increasingly selfish behaviour as the payoff matrix is scaled down. In contrast, Mistral Large demonstrates the opposite tendency, with its behaviour shifting counter to the trend observed in GPT-4o.
The performance trajectories of the Claude 3.5 Haiku model in Figure 3 exhibit a general downward trend as the number of iterations increases. The very low trajectory, predominantly situated in the upper quadrant (favouring Option A), displays a gradual shift towards the lower quadrant (Option B), interrupted by a minor peak at Round 5 that suggests a notable degree of strategic volatility. Conversely, the high and ordinary trajectories are primarily concentrated in the lower domain, indicating a cooperative tendency designed to yield a higher average expected utility. The very low line demonstrates a significantly greater inclination towards the upper spectrum compared to the high line, a divergence that may be partly attributed to linguistic bias [bang2023multitask]. Specifically, within the context of the Vietnamese language combined with a payoff scalar of $\lambda=0.1$, Claude 3.5 Haiku appears to misinterpret the unit magnitude of the payoff matrix, resulting in the erroneous maximization of penalties rather than their intended minimization.
@@
Regarding the Mistral Large model, empirical observations indicate that higher payoff values correlate with an increased likelihood of defection, a behaviour consistent with the dominant strategy concept in game theory. Furthermore, a distinct visual correlation is evident, characterized by a substantial spike at Round 2. This anomaly suggests a retaliatory mechanism triggered by a defection in the initial round, or alternatively [akata2025], represents a probing tactic aimed at minimizing the maximum potential penalty.
4.2 LLM behaviours in Public Goods Game
We now turn to the multi-agent Public Goods Game to address RQ2: how do LLM agents behave in a collective group social dilemma, and how does this behaviour depend on incentives, time, and framing (model, personality, language).